3/4x-5=4
Please help me figure out this problem, I have been stuck on it for awhile.
this is what i've done but I do not have a clue if its right.
3(3/4x)-5=4
3x-5=4
3*3-5=4
9-5=4
x=3

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

(3/4)X-5=4 OR (3/4)X=4+5 OR (3/4)X=9 OR 3X=36 OR X=36/3 OR X=12
PROOF (3/4)12-5=4 OR 9-5=4 OR 4=4

Answer by stanbon(75887)

(3/4)x-5=4
Add 5 to both sides to get:
(3/4)x=9
Multiply both sides by (4/3) to get:
x=9(4/3)
Simplify, as follows:
x=12

Cheers,
Stan H.